# FeNNol: an Efficient and Flexible Library for Building Force-field-enhanced Neural Network Potentials

### v1

@@ V.4 Native MD engine
FeNNol provides an efficient native engine to perform molecular dynamics simulations. The engine is written purely in Jax, is completely GPU-resident and leverages Jax’s asynchronous dispatch capabilities for efficient dynamics even for small systems. Multiple thermostats are available (Langevin, Bussi-Parinello or Nose-Hoover) and the integration scheme used is based on symplectic time-splitting methods and is adapted depending on the thermostat (for example it reverts to the velocity Verlet integrator for NVE MD and to the BAOAB integrator Leimkuhler and Matthews (2013) for Langevin MD). Simulations can be performed either in vacuum or with periodic boundary conditions (with interactions with all images or minimum image convention). A current limitation of the engine, however, is that constant pressure simulations are not yet implemented. Furthermore, multi-GPU parallelization is not yet available.
@@ V.5 Performance benchmarks
In the remaining of this section, we provide a brief study of simulation accuracy with respect to Jax matrix multiplication level of precision (via energy conservation tests) and performance benchmarks for our implementation of the ANI-2x Devereux et al. (2020) model. The model is implemented only with standard FeNNol modules (i.e. without custom CUDA kernels). Parameters for the neural network are obtained from the TorchANI Gao et al. (2020) implementation and converted to Flax variables. We only use the first from the ensemble of eight neural network that compose the ANI-2x model. All our tests were performed with FeNNol’s native MD engine using Jax 0.4.25 with jaxlib 0.4.23 and CUDA 11.8 on a NVIDIA RTX 3090 GPU.
For our accuracy tests, we compare single (float32) and double precision (float64) representations. For single precision, we also compare the various matrix multiplications precisions provided by Jax that are denoted as "default" (which uses bfloat16 numbers), "high" (which uses tensorfloat32) and "highest" (which uses full float32). Figure 3 shows the total energy of a periodic box of 216 water molecules over a 1ns NVE simulation with a timestep of 0.1 fs (this small timestep was chosen to minimize integration error and focus on precision errors) with the different precision configurations. All the simulations start from the same off-equilibrium configuration and velocities are initialized to zero (this results in an average temperature of 315K with a standard deviation of 12K). It is clear from the figure, that for this simulation setup, both "default" and "high" configurations are not precise enough to obtain good energy conservation. Furthermore, we see a non-negligible shift in the initial energy compared to float64. The "highest" float32 configuration provides acceptable energy conservation (fluctuations below 0.1% of the initial energy, as indicated by the green box in the figure) and no initial energy shift compared to the full float64 simulation. From a performance perspective, differences between the three single precision configurations are small, while double precision is significantly more expensive on this GPU (about a factor of 4). Note that this difference between single and double precision is usually less important with HPC GPUs. For example using double precision on a A100 GPU only adds a 15% overhead. Given these tests, we set the default precision for FeNNol’s MD engine to "highest" float32 and warn users of Jax that the default precision settings might not be sufficient in the context of MD simulations.
@@
Figure 4: Performance (in ns simulated per day) for various system sizes. The black dashed line corresponds to the AMOEBA force field simulated using Tinker-HP. The solid lines correspond to ANI-2x with FeNNol’s implementation and different MD engines. The dotted lines correspond to ANI-2x with TorchANI’s implementation. All simulations were performed on a single NVIDIA RTX 3090 GPU using float32 precision, a 0.5fs timestep and a Langevin thermostat set at 300K.
@@
We now test the performance of our implementation on four water boxes which we denote "watersmall" (216 molecules), "waterbox" (500 molecules), "waterbig" (1600 molecules) and "waterhuge" (4000 molecules), and on the solvated DHFR (around 20k atoms). We compare our implementation to the base TorchANI implementation (with ASE as MD engine). We also compare the performance of ANI-2x to the AMOEBA polarizable force field Ponder et al. (2010) simulated using Tinker-HP. All simulations were performed on a single NVIDIA RTX 3090 GPU using float32 precision, a 0.5fs timestep, a Langevin thermostat set at 300K. Both FeNNol and TorchANI are set to only use the first out of the eight neural networks that compose the ANI-2x model. Figure 4 shows the simulation performance (in ns simulated per day) for the different setups. In order to obtain fairer comparisons, we implemented the minimum image convention in TorchANI, which already provides a significant increase in performance for these systems. Note that we could not simulate the larger systems with TorchANI due to allocations exceeding our GPU’s memory. If we first compare TorchANI to FeNNol both running with the ASE MD engine, we see that FeNNol gains almost a factor of two in performance which is mainly due to the efficient JIT compilation of Jax 111a fairer comparison could be achieved by using the recently introduced JIT compilation tools in Pytorch 2 but that would require a complete rewrite of TorchANI with fixed array shapes.. Since ASE integrates the equations of motion on the CPU, using this engine requires frequent data transfers between CPU and GPU. Using FeNNol’s native MD engine, we alleviate this issue with a fully GPU-resident and asynchronous simulation, which considerably increase performance for smaller systems (nearly a factor of three for the smallest system). For larger systems, the main bottleneck becomes the construction of neighbour lists, which in the current implementation, scales quadratically with system size (compared to the linear scaling of the ANI-2x model itself). The impact of this quadratic scaling can be pushed back to larger system sizes by using a neighbour list "skin" and to reconstruct the full neighbour list only once every few steps. If we add a buffer of 2 Å and, following the rule used in Tinker-HP, reconstruct the neighbour list once every 40 fs (80 steps), we obtain a significant improvement, reaching performance close to the AMOEBA force field on the smaller systems. This performance increase is due to both a gain in neighbour list construction and a better leveraging of Jax’s asynchronous dispatch when using the additional buffer. Note that this comparison only takes into account the raw performance of the models with a fixed integration scheme. Typical MD runs with force fields however can usually afford larger timesteps and use advanced integration schemes (for example multi-timesteppingLagardere, Aviat, and Piquemal (2019)) which are not compatible with current ML potentials and will require architectural evolutions.
@@
The results reported here were measured with a RTX 3090 GPU but we found very similar performance for a A100 HPC GPU (with a slight increase in performance for the larger systems). For both GPUs, utilization is constantly close to 100% for all systems except the smallest "watersmall" system (with the FeNNol native engine). For the A100 GPU, power usage reaches 70% of the cap value (280W / 400W) for the larger system. For the RTX 3090, power usage stays close to the cap value (350W) for all systems except the smallest one, showing nearly maximal GPU occupancy. We can thus expect performance gains with the newer generation RTX 4090 that are close to theoretical values (i.e. at least a factor of two).
@@
This brief study shows that neural network potentials can indeed reach performance close to force fields (at least the more expensive and accurate polarizable ones) and that the efficiency of the model implementation and the simulation engine are both critical.
@@ VI.3 Learning distributed multipoles
For the last example, we trained an equivariant model to reproduce distributed MB-ISA multipoles (up to quadrupoles) from the Pubchem subset of the SPICE dataset Eastman et al. (2023). We selected a CRATE embedding with radial, angular and equivariant resources with three interaction layers of cutoff 3.5Å. We used 32 equivariant channels with maximum degree $l_{\text{max}}=2$ and perform one tensor product per layer. We obtain the atomic dipoles and quadrupoles from a linear combination of the equivariant channels of the corresponding irreducible representation. To obtain total-charge-constrained partial charges, we use the scalar embedding produced by CRATE to define a weight on each atom that is then used to distribute all the valence electrons in the system (included ionic charges). This constrains the total number of electrons and prevents the atoms from unphysically depleting more than their number of valence electrons with too large positive charges. We kept 10% of the 1.4 million configurations as a validation set. Furthermore, we preprocessed the quadrupoles in the dataset to make them traceless. We used a similar training setup as in the previous two examples and applied the same weight to all the multipole components in the loss function. Training completed in 13.5 hours and the model achieved a MAE of 0.004 $e$ on charges, 0.0017 $e$Å on dipole components and 0.0012 $e$Å2 on quadrupole components on the validation set. For reference, the MAE for a trivial model that predicts only neutral spherical atoms is 0.27 $e$ for charges, 0.019 $e$Å  for dipoles and $0.0087$ $e$Å2 for quadrupoles.
@@ VII Conclusion
Finally, we provide three main approaches for running MD simulations with ML models: an ASE calculator, an interface to Tinker-HP via Deep-HP Inizan et al. (2022) and an efficient native MD engine. The ASE calculator allows to transparently use all of the functionalities and benefit from the flexibility of the Atomic Simulation Environment Larsen et al. (2017). Similarly, the Deep-HP interface enables the use of all of Tinker-HP’s advanced functionalities and its massive parallelization capabilities. Furthermore it allows for mixed ML/MM simulations as shown in ref. Inizan et al., 2022. Finally, the native MD engine allows to perform efficient MD simulations within the FeNNol package by leveraging the full capabilities of Jax’s JIT compilation and asynchronous dispatch. We showed that our implementation of the ANI-2x model can reach simulation speeds on par with an optimized version of the AMOEBA polarizable force field.
@@ Appendix A Detailed description of the CRATE embedding
where $C^{\lambda_{1}\lambda_{2}\lambda}_{m_{1}m_{2}m}$ are Clebsch–Gordan coefficients for the basis of real spherical harmonics, the index $p$ enumerates tensor product paths that produce a certain irrep $l$ (the user can choose to include either all the paths that produce a certain $l$ or only the ones which conserve the parity of the destination irrep). The trainable parameters $W_{\text{\tiny{path}}}^{(l,t)}$ combine the different tensor product paths to conserve the number of irreps. This is similar to the scheme proposed in ref. Kozinsky et al., 2023 and is key to prevent the explosion of the number of paths after iterating tensor products. Equation (14) corresponds to an update of $V_{i}$ followed by channel mixing. We initialize as $\hat{V}_{i}^{(l,0)}=\hat{V}_{i}^{(l-1)}$ and keep the last iteration as the current equivariant embedding $\hat{V}_{i}^{(l)}=\hat{V}_{i}^{(l,N_{\text{\tiny{TP}}})}$.

### v2

@@ V.4 Native MD engine
FeNNol provides an efficient native engine to perform molecular dynamics simulations, which can be invoked with the fennol_md command. The engine is written purely in Jax, is completely GPU-resident and leverages Jax’s asynchronous dispatch capabilities for efficient dynamics even for small systems. Multiple thermostats are available (Langevin, Bussi-Parinello or Nose-Hoover) and the integration scheme used is based on symplectic time-splitting methods and is adapted depending on the thermostat (for example it reverts to the velocity Verlet integrator for NVE MD and to the BAOAB integrator Leimkuhler and Matthews (2013) for Langevin MD). Simulations can be performed either in vacuum or with periodic boundary conditions (with interactions with all images or minimum image convention). A current limitation of the engine, however, is that constant pressure simulations are not yet implemented. Furthermore, multi-GPU parallelization is not yet available.
@@ V.5 Performance benchmarks
In the remaining of this section, we provide a brief study of simulation accuracy with respect to Jax matrix multiplication level of precision (via energy conservation tests) and performance benchmarks for our implementation of the ANI-2x Devereux et al. (2020) model. The model is implemented only with standard FeNNol modules (i.e. without custom CUDA kernels). Parameters for the neural network are obtained from the TorchANI Gao et al. (2020) implementation and converted to Flax variables. We only use the first from the ensemble of eight neural networks that compose the ANI-2x model. All our tests were performed with Jax 0.4.25 with jaxlib 0.4.23 and CUDA 11.8. We used either a NVIDIA RTX 3090 or A100 GPU with respectively 24GB and 40GB of memory.
For our accuracy tests, we compare single (float32) and double precision (float64) representations. For single precision, we also compare the various matrix multiplications precisions provided by Jax that are denoted as "default" (which uses bfloat16 numbers), "high" (which uses tensorfloat32) and "highest" (which uses full float32). Figure 3 shows the total energy of a periodic box of 216 water molecules over a 1 ns NVE simulation with a timestep of 0.1 fs (this small timestep was chosen to minimize integration error and focus on precision errors) with the different precision configurations. All the simulations start from the same off-equilibrium configuration and velocities are initialized to zero (this results in an average temperature of 315K with a standard deviation of 12K). It is clear from the figure, that for this simulation setup, both "default" and "high" configurations are not precise enough to obtain good energy conservation. Furthermore, we see a non-negligible shift in the initial energy compared to float64. The "highest" float32 configuration provides acceptable energy conservation (fluctuations below 0.1% of the initial energy, as indicated by the green box in the figure) and no initial energy shift compared to the full float64 simulation. From a performance perspective, differences between the three single precision configurations are small, while double precision is significantly more expensive on the RTX 3090 GPU (about a factor of 4). Note that this difference between single and double precision is less important with HPC GPUs. For example using double precision on a A100 GPU only adds a 15% overhead. Given these tests, we set the default precision for FeNNol’s MD engine to "highest" float32 and warn users of Jax that the default precision settings might not be sufficient in the context of MD simulations.
@@
Figure 4: Performance (in ns simulated per day) for various system sizes. The black dashed line corresponds to the AMOEBA force field simulated using Tinker-HP. The solid lines correspond to ANI-2x with FeNNol’s implementation and different MD engines. The dotted lines correspond to ANI-2x with TorchANI’s implementation (min. img.=minimum image convention). All simulations use float32 precision, a 0.5fs timestep and a Langevin thermostat set at 300K. Simulations with TorchANI and FeNNol’s native MD engine were performed on a RTX 3090 GPU. Simulations with Deep-HP were performed on a A100 GPU.
FeNNol FeNNol+skin FeNNol+ASE FeNNol+Deep-HP TorchANI TorchANI+Deep-HP
model implementation Jax Jax Jax Jax Pytorch Pytorch
neighbour list ($\order{N^{2}}$ ) ($\order{N^{2}}/N_{\text{update}}$ ($\order{N^{2}}$ ) ($\order{N}$ ) ($\order{N^{2}}$ ) ($\order{N}$ )
MD Engine FeNNol(Jax) FeNNol(Jax) ASE(Numpy) Tinker-HP(CUDA) ASE(Numpy) Tinker-HP(CUDA)
Table 1: Summary of the different implementations that are compared in the performance benchmark of section V.5. The "neighbour list" line shows the complexity of the neighbour list construction. The column headers refer to the legend in Figure 4.
We now test the performance of our implementation on water boxes of varying sizes (from 648 atoms to 500K atoms) and on the solvated DHFR (around 20K atoms). We compare our implementation (with ASE, Deep-HP or our native MD engine) to the base TorchANI implementation (with ASE or Deep-HP as MD engine). We also compare the performance of ANI-2x to the AMOEBA polarizable force field Ponder et al. (2010) simulated using Tinker-HP. Table 1 summarizes the differences between the implementations that we compare in this section. All simulations use float32 precision, a 0.5 fs timestep, a Langevin thermostat set at 300K and periodic boundary conditions. Both FeNNol and TorchANI are set to only use the first out of the eight neural networks that compose the ANI-2x model. Simulations with TorchANI and FeNNol’s native MD engine were performed on a RTX 3090 GPU with 24GB of memory. Simulations with FeNNol+Deep-HP were performed on a A100 GPU with 40GB of memory.
@@
Figure 4 shows the simulation performance (in ns simulated per day) for the different setups. In order to obtain fairer comparisons, we implemented the minimum image convention in TorchANI, which already provides a significant increase in performance for these systems. Note that we could not simulate the larger systems with TorchANI due to allocations exceeding our GPU’s memory. If we first compare TorchANI to FeNNol both running with the ASE MD engine, we see that FeNNol gains almost a factor of two in performance which is mainly due to the efficient JIT compilation of Jax (note that a more direct comparison could be achieved by using the recently introduced JIT compilation tools in Pytorch 2 but that would require a complete rewrite of TorchANI with fixed array shapes). Since ASE integrates the equations of motion on the CPU, using this engine requires frequent data transfers between CPU and GPU. Using FeNNol’s native MD engine, we alleviate this issue with a fully GPU-resident and asynchronous simulation, which considerably increase performance for smaller systems (nearly a factor of three for the smallest system). For larger systems, the main bottleneck becomes the construction of neighbour lists, which in the current implementation, scales quadratically with system size (compared to the linear scaling of the ANI-2x model itself). The impact of this quadratic scaling can be pushed back to larger system sizes by using a neighbour list "skin" and to reconstruct the full neighbour list only once every few steps. If we add a buffer of 2 Å and, following the rule used in Tinker-HP, reconstruct the neighbour list once every 40 fs (80 steps), we obtain a significant improvement, reaching performance close to the AMOEBA force field on the smaller systems. This performance increase is due to both a gain in neighbour list construction and a better leveraging of Jax’s asynchronous dispatch when using the additional buffer. Note that this comparison only takes into account the raw performance of the models with a fixed integration scheme. Typical MD runs with force fields however can usually afford larger timesteps and use advanced integration schemes (for example multi-timesteppingLagardere, Aviat, and Piquemal (2019)) which are not compatible with current ML potentials and will require architectural evolutions.
For systems of more than 20K atoms, the quadratic scaling of neighbour lists construction in FeNNol’s native MD engine takes a significant toll on performance. We thus switched to Deep-HP – which leverages the efficient cell lists implemented in Tinker-HP – for systems between 20K and 500K atoms. For these large systems, we see similar scaling trends with FeNNol+Deep-HP compared to AMOEBA with about a factor of 3 in raw performance. Note that the FeNNol+Deep-HP interface is in very early development, so that we expect this gap to shrink with further optimizations and a tighter integration between both codes. In particular, we expect that performing all the model preprocessing (which does not require automatic differentiation) inside Tinker-HP will provide a boost in performance and significant memory savings. When comparing to the raw performance of TorchANI+Deep-HP, we see again a factor $\sim$2.5 with respect to FeNNol.
@@
The results reported here were measured with a RTX 3090 GPU for smaller systems and a A100 HPC GPU for larger systems. We also tested the A100 GPU on the smaller systems and did not see significant differences with respect to the RTX GPU. For both GPUs, utilization is constantly close to 100% for all systems except the smallest one. For the A100 GPU, power usage reaches 70% of the cap value (280W / 400W) for the larger systems. For the RTX 3090, power usage stays close to the cap value (350W) for all systems except the smallest one, showing nearly maximal GPU occupancy. We can thus expect performance gains with the newer generation RTX 4090 that are close to theoretical values (i.e. at least a factor of two).
@@
This brief study shows that neural network potentials can indeed reach performance close to force fields (at least the more expensive and accurate polarizable ones) and that the efficiency of the model implementation and the MD engine are both critical.
@@ VI.3 Learning distributed multipoles
For the last example, we trained an equivariant model to reproduce distributed MB-ISA multipoles (up to quadrupoles) from the Pubchem subset of the SPICE dataset Eastman et al. (2023). We selected a CRATE embedding with radial, angular and equivariant resources with three interaction layers of cutoff 3.5Å. We used 32 equivariant channels with maximum degree $l_{\text{max}}=2$ and perform one tensor product per layer. We obtain the atomic dipoles and quadrupoles from a linear combination of the equivariant channels of the corresponding irreducible representation. To obtain total-charge-constrained partial charges, we use the scalar embedding produced by CRATE to define a weight on each atom that is then used to distribute all the valence electrons in the system (including eventual ionic charges). This constrains the total number of electrons and prevents the atoms from unphysically depleting more than their number of valence electrons with too large positive charges. We kept 10% of the 1.4 million configurations as a validation set. Furthermore, we preprocessed the quadrupoles in the dataset to make them traceless. We used a similar training setup as in the previous two examples and applied the same weight to all the multipole components in the loss function. Training completed in 14 hours and the model achieved a MAE of 0.004 $e$ on charges, 0.0017 $e$Å on dipole components and 0.0013 $e$Å2 on quadrupole components on the validation set. For reference, the MAE for a trivial model that predicts only neutral spherical atoms is 0.27 $e$ for charges, 0.019 $e$Å  for dipoles and $0.0087$ $e$Å2 for quadrupoles.
@@ VII Conclusion
Finally, we provide three main approaches for running MD simulations with ML models: an ASE calculator, an interface to Tinker-HP via Deep-HP Jaffrelot Inizan et al. (2023) and an efficient native MD engine. The ASE calculator allows to transparently use all of the functionalities and benefit from the flexibility of the Atomic Simulation Environment Larsen et al. (2017). Similarly, the Deep-HP interface enables the use of all of Tinker-HP’s advanced functionalities and its massive parallelization capabilities. Furthermore it allows for mixed ML/MM simulations as shown in ref. Jaffrelot Inizan et al., 2023. We showed that our early verision of the interface enables simulations of systems up to 500K atoms on a single A100 GPU with performance and scaling comparable to the AMOEBA force fields. Finally, FeNNol’s native MD engine allows to perform very efficient MD simulations of intermediate size systems (more than 20K atoms) directly within the FeNNol package by leveraging the full capabilities of Jax’s JIT compilation and asynchronous dispatch. We showed that our implementation of the ANI-2x model can reach simulation speeds nearly on par with an optimized version of the AMOEBA polarizable force field.
@@ Appendix A Detailed description of the CRATE embedding
where $C^{\lambda_{1}\lambda_{2}\lambda}_{m_{1}m_{2}m}$ are Clebsch–Gordan coefficients for the basis of real spherical harmonics, the index $p$ enumerates tensor product paths that produce a certain irrep of degree $\lambda$ (the user can choose to include either all the paths that produce a certain $\lambda$ or only the ones which conserve the parity of the output irrep). The trainable parameters $W_{\text{\tiny{path}}}^{(l,t)}$ combine the different tensor product paths to conserve the number of irreps. This is similar to the scheme proposed in ref. Kozinsky et al., 2023 and is key to prevent the explosion of the number of paths after iterating tensor products. Equation (14) corresponds to an update of $V_{i}$ followed by channel mixing. We initialize as $\hat{V}_{i}^{(l,0)}=\hat{V}_{i}^{(l-1)}$ and keep the last iteration as the current equivariant embedding $\hat{V}_{i}^{(l)}=\hat{V}_{i}^{(l,N_{\text{\tiny{TP}}})}$.